# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ D.2 Baselines and Hyper-parameter Settings
Method Average Reward Average Cost Safety Rate Inference Time (s)
Alpaca-7B Base 6.15 ($\pm$ 1.51) 1.33 29.47% 1.1
RECONTROL 6.2 ($\pm$ 1.56) 1.33 29.5% 1.7
RECONTROL + Lagrangian multiplier 6.19 ($\pm$ 1.50) 1.33 29.7% 2
Best-of-N + Lagrangian multiplier $N=100$, $\lambda=5$ 5.35 ($\pm$ 1.62) -0.46 48.22% 29
Best-of-N + Lagrangian multiplier $N=200$, $\lambda=5$ 5.25 ($\pm$ 1.64) -0.72 54.2% 58
Best-of-N + Lagrangian multiplier $N=500$, $\lambda=5$ 6.04 ($\pm$ 1.85) -1.27 52.17% 145
Best-of-N + Lagrangian multiplier $N=500$, $\lambda=10$ 5.51 ($\pm$ 1.66) -1.44 54.01% 145
Best-of-N + Augmented safety $N=200$ 7.51 ($\pm$ 1.89) 0.67 60.07% 58
Best-of-N + Augmented safety $N=500$ 7.78 ($\pm$ 2.09) 0.42 65.74 % 145
Beam search + Lagrangian multiplier $N=128$, $\lambda=5$ 6.58 ($\pm$ 1.95) -1.02 50.19% 32
Beam search + Lagrangian multiplier $N=256$, $\lambda=5$ 6.69 ($\pm$ 2.08) -1.28 52.43% 60
Beam search + Augmented safety $N=128$ 8.29 ($\pm$ 2.02) 0.64 58.89% 39
Beam search + Augmented safety $N=256$ 8.69 ($\pm$ 2.15) 0.55 61.79 % 82
ARGS $\omega=2.5$ 6.74 ($\pm$ 1.70) 1.47 28.19% 82
ARGS + Lagrangian multiplier $\omega=2.5$ 3.21 ($\pm$ 1.59) -0.85 75.8% 111
ARGS + Cost Model $\omega=2.5$ 0.19 ($\pm$ 1.65) -2.21 81.6% 78
InferenceGuard (N=128) 7.08 ($\pm$ 2.49) -0.63 88.14% 65
[…]
Beaver-7B-v3 Base 5.83 ($\pm$ 1.62) -2.89 75.89% 1.2
RECONTROL 5.9 ($\pm$ 1.56) -2.90 75.9% 2
RECONTROL + Lagrangian multiplier 5.91 ($\pm$ 1.50) -2.91 75.9% 2.6
Best-of-N + Lagrangian multiplier $N=100$ 6.52 ($\pm$ 1.88) -3.63 85.7% 40
Best-of-N + Lagrangian multiplier $N=200$ 6.61 ($\pm$ 1.89) -3.62 85.8% 58
Best-of-N + Augmented safety $N=100$ 8.55 ($\pm$ 1.58) -2.96 97.23% 40
Best-of-N + Augmented safety $N=200$ 9.01 ($\pm$ 1.63) -2.98 97.76% 58
Beam search + Lagrangian multiplier $N=64$, $\lambda=5$ 8.33 ($\pm$ 1.79) -4.09 87.08% 36
Beam search + Lagrangian multiplier $N=128$, $\lambda=5$ 8.63 ($\pm$ 1.80) -4.21 87.35 % 64
Beam search + Augmented safety $N=64$ 9.84 ($\pm$ 1.4) -2.93 95.38 % 22
Beam search + Augmented safety $N=128$ 10.31 ($\pm$ 1.37) -2.94 97.36% 39
ARGS $\omega=2.5$ 6.72 ($\pm$ 1.83) -2.59 78.5% 94
ARGS $\omega=2.5$ + Lagrangian multiplier 2.26 ($\pm$ 1.56) -1.64 81% 127
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 1.37) -3.27 98.4% 90
InferenceGuard (N=128) 10.26 ($\pm$ 1.42) -2.96 99.7% 39
InferenceGuard with Critic (N=128) 10.27 ($\pm$ 1.50) -2.94 100% 39
D.3 Critic Network and Training Process

### v3

@@ D.2 Experiment Settings and Results
Dataset Method Average Reward Safety Rate Inference Time (s)
Vicuna-7B HEx-PHI Base 4.69 ($\pm$ 1.36) 48% 1.2
RE-Control 4.75 ($\pm$ 1.31) 49.33% 1.58
RE-Control + Lagrangian multiplier 4.65 ($\pm$ 1.33) 50.7% 1.75
Best-of-N + Lagrangian multiplier 5.22 ($\pm$ 1.39) 79.3% 9.08
Best-of-N + Augmented safety 6.46 ($\pm$ 1.51) 92.6% 10.04
Beam search + Lagrangian multiplier 5.70 ($\pm$ 1.57) 83% 7.2
Beam search + Augmented safety 7.57 ($\pm$ 1.67) 89.33% 11.59
ARGS $\omega=2.5$ 5.67 ($\pm$ 1.45) 47% 68.23
ARGS $\omega=2.5$ + Lagrangian multiplier 1.72 ($\pm$ 1.96) 93.33% 79.28
ARGS $\omega=2.5$ + Cost Model 0.07 ($\pm$ 1.60) 96% 69.38
InferenceGuard 6.90 ($\pm$ 2.08) 96.67% 11.01
InferenceGuard with Critic 6.99 ($\pm$ 2.1) 96.67% 13.29
Vicuna-7B HH-RLHF Base 5.82 ($\pm$ 1.56) 95% 1.17
RE-Control 5.9 ($\pm$ 1.55) 95.13% 1.45
RE-Control + Lagrangian multiplier 5.85 ($\pm$ 1.50) 95.4% 2.09
Best-of-N + Lagrangian multiplier 6.97 ($\pm$ 2.54) 97.24% 8.32
Best-of-N + Augmented safety 8.33 ($\pm$ 1.95) 98.36% 8.58
Beam search + Lagrangian multiplier 8.05 ($\pm$ 2.25) 97.54% 11.28
Beam search + Augmented safety 9.61 ($\pm$ 2.10) 98.23% 11.84
ARGS $\omega=2.5$ 6.83 ($\pm$ 1.83) 96.2% 70.4
ARGS $\omega=2.5$ + Lagrangian multiplier 2.02 ($\pm$ 1.79 ) 97.54% 73.74
ARGS $\omega=2.5$ + Cost Model 0.46 ($\pm$ 1.73 ) 98.96% 72
InferenceGuard 9.49 ($\pm$ 2.16) 98.97% 11.47
InferenceGuard with Critic 9.48 ($\pm$ 2.16) 98.95% 11.73
Table 4: Performance Comparison using Llama-3.18B-Instruct, evaluated by reward model QRM-Llama-3.1-8B, on Datasets PKU-SafeRLHF and HH-RLHF using $N=128$, $\lambda=5$
Dataset Method Average Reward Safety Rate Inference Time (s)
Llama3.1-8B-Instruct PKU-SafeRLHF Base 0.53 ($\pm$ 0.12) 53.22% 1.13
RE-Control 0.55 ($\pm$ 0.22) 52.7% 1.52
RE-Control + Lagrangian multiplier 0.51 ($\pm$ 0.21) 54.01% 2.11
Best-of-N + Lagrangian multiplier 0.35 ($\pm$ 0.17) 81.29% 5.70
Best-of-N + Augmented safety 0.47 ($\pm$ 0.16) 80.89% 5.71
Beam search + Lagrangian multiplier 0.39 ($\pm$ 0.16) 81.69% 8.60
Beam search + Augmented safety 0.46 ($\pm$ 0.16) 81.43% 8.68
ARGS $\omega=2.5$ 0.55 ($\pm$ 0.21) 52.6% 74.53
ARGS $\omega=2.5$ + Lagrangian multiplier 0.25 ($\pm$ 0.23 ) 73.33% 78.22
ARGS $\omega=2.5$ + Cost Model 0.05 ($\pm$ 0.29) 94.0% 71.61
InferenceGuard 0.44 ($\pm$ 0.13) 90.91% 10.69
Llama3.1-8B-Instruct HH-RLHF Base 0.09 ($\pm$ 0.27) 23.5% 0.82
RE-Control 0.11 ($\pm$ 0.29) 23.3% 1.07
RE-Control + Lagrangian multiplier 0.10 ($\pm$ 0.21) 24.7% 1.26
Best-of-N + Lagrangian multiplier 0.31 ($\pm$ 0.21) 82.37% 5.53
Best-of-N + Augmented safety 0.42 ($\pm$ 0.19) 84.5% 5.61
Beam search + Lagrangian multiplier 0.31 ($\pm$ 0.21) 83.96% 7.12
Beam search + Augmented safety 0.42 ($\pm$ 0.2) 83.15% 7.08
ARGS $\omega=2.5$ 0.52 ($\pm$ 0.29 ) 49.8% 68.9
ARGS $\omega=2.5$ + Lagrangian multiplier 0.24 ($\pm$ 0.19) 87.8% 71.4
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 0.29 ) 96.01% 64.2
InferenceGuard 0.38 ($\pm$ 0.14) 98.45% 8.11
D.3 Ablation Study
Effect of $\lambda$ and sample size $N$ on baselines
To further evaluate the safety-performance trade-off in baseline methods, we compare decoding strategies across a range of Lagrangian multipliers $\lambda$ and sample sizes $N$ for Best-of-N, Beam Search, and ARGS, under varying $\lambda$ values and sample budgets, as shown in Table 5 for Alpaca-7B and Beaver-7B, and Table 6 for Vicuna-7B. For Beam Search specifically, we apply an alternative technique to enhance sampling diversity by reducing the frequency with which $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ falls below a threshold; we refer to this variant as Beam Search Lag. + Freq. While safety rates improve as $\lambda$ increases, task reward often degrades sharply, particularly for ARGS. In contrast, InferenceGuard achieves a more favorable balance without requiring fine-tuning of $\lambda$.
Table 5: Performance Comparison of Lagrangian Multiplier-Based Methods on Dataset PKU-SafeRLHF using Different $\lambda$ and $N$
Method $\lambda$ Average Reward Safety Rate Inference Time (s)
Alpaca-7B InferenceGuard $N=128$ 7.08 ($\pm$ 2.49) 88.14% 16.25
Best-of-N Lag. $N=128$ 0 7.92 ($\pm$ 1.43) 34.12% 8.03
1 7.42 ($\pm$ 1.72) 42.82% 7.95
2.5 6.64 ($\pm$ 1.89) 48.75% 8.89
5 5.97 ($\pm$ 1.82) 51.25% 9.62
10 5.54 ($\pm$ 1.64) 52.70% 8.88
Best-of-N Lag. $N=256$ 0 8.41 ($\pm$ 1.45) 33.86% 14.53
1 7.82 ($\pm$ 1.75) 42.95% 13.73
2.5 6.78 ($\pm$ 2.01) 51.25% 13.78
5 6.04 ($\pm$ 1.85) 54.41% 13.15
10 5.51 ($\pm$ 1.69) 55.20% 14.98
Beam Search Lag. $N=128$ 0 8.90 ($\pm$ 1.71) 27.80% 8.72
1 8.17 ($\pm$ 2.10) 41.37% 8.75
2.5 7.37 ($\pm$ 2.22) 50.46% 8.77
5 6.58 ($\pm$ 1.95) 50.19% 8.01
10 5.85 ($\pm$ 1.79) 49.93% 8.69
Beam Search Lag. + Freq. $N=128$ 1 8.15 ($\pm$ 2.09) 43.70% 8.66
2.5 7.42 ($\pm$ 2.21) 50.49% 8.74
5 6.59 ($\pm$ 1.98) 50.6% 8.42
10 5.85 ($\pm$ 1.79) 49.94% 9.55
Beam Search Lag. $N=256$ 0 9.35 ($\pm$ 1.83) 30.43% 14.31
1 8.51 ($\pm$ 2.17) 43.21% 14.63
2.5 7.58 ($\pm$ 2.25) 50.46% 14.65
5 6.69 ($\pm$ 2.08) 52.43% 15.03
10 5.90 ($\pm$ 1.82) 53.10% 14.76
Beam Search Lag. + Freq. $N=256$ 1 8.56($\pm$ 2.17) 45.22% 14.7
2.5 7.61 ($\pm$ 2.29) 50.93% 14.68
5 6.66 ($\pm$ 2.12) 52.9% 13.14
10 5.89 ($\pm$ 1.84) 52.25% 13.09
ARGS Lag. 0 6.74 ($\pm$ 1.70) 28.19% 58.58
1 4.07 ($\pm$ 1.64) 65.6% 62.28
2.5 3.98 ($\pm$ 1.61) 66.0% 69.72
5 3.21 ($\pm$ 1.59) 75.8% 63.42
10 1.23 ($\pm$ 1.63) 79.2% 61.14
Beaver-7B-v3 InferenceGuard. $N=128$ 10.26 ($\pm$ 1.42) 99.7% 9.75
Best-of-N Lag.$N=64$ 0 8.68 ($\pm$ 1.37) 77.07% 4.52
1 8.47 ($\pm$ 1.45) 81.69% 4.43
2.5 7.95 ($\pm$ 1.64) 85.11% 4.88
5 7.06 ($\pm$ 1.77) 87.48% 4.67
10 6.22 ($\pm$ 1.69) 88.14% 4.43
Best-of-N Lag.$N=128$ 0 9.15 ($\pm$ 1.32) 76.82% 7.86
1 8.92 ($\pm$ 1.43) 81.69% 8.02
2.5 8.35 ($\pm$ 1.64) 84.19% 7.72
5 7.30 ($\pm$ 1.80) 87.20% 7.92
10 6.31 ($\pm$ 1.76) 87.62% 7.82
Beam Search Lag. $N=64$ 0 11.02 ($\pm$ 1.34) 74.70% 5.18
1 10.64 ($\pm$ 1.37) 82.35% 5.9
2.5 9.99 ($\pm$ 1.58) 87.62% 5.01
5 9.84 ($\pm$ 1.4) 95.38 % 5.54
10 7.60 ($\pm$ 1.82) 89.20% 4.97
Beam Search Lag. $N=128$ 0 10.54 ($\pm$ 1.29) 74.44% 8.92
1 10.25 ($\pm$ 1.41) 82.74% 9.05
2.5 9.57 ($\pm$ 1.60) 87.10% 9.6
5 10.31 ($\pm$ 1.37) 97.36% 9.75
10 7.34 ($\pm$ 1.86) 88.41% 9.9
ARGS Lag. 0 6.72 ($\pm$ 1.83) 78.5% 67.15
1 4.01 ($\pm$ 1.61) 80.9% 75.0
2.5 3.67 ($\pm$ 1.61) 80.65% 74.5
5 2.26 ($\pm$ 1.56) 81% 73.74
10 0.95 ($\pm$ 1.67) 90.8% 68.75
Table 6: Performance Comparison of Lagrangian Multiplier-Based Methods on Datasets HEx-PHI and HH-RLHF using Different $\lambda$ and $N$ (Vicuna-7B-v1.5)
Dataset Method $\lambda$ Average Reward Safety Rate Inference Time (s)
HEx-PHI InferenceGuard. $N=128$ 6.90 ($\pm$ 2.08) 96.67% 11.01
Best-of-N Lag. $N=64$ 0 6.97 ($\pm$ 1.27) 34.33% 6.5
1 6.12 ($\pm$ 1.52) 66% 6.4
2.5 5.54 ($\pm$ 1.55) 74% 6.85
5 4.86 ($\pm$ 1.52) 78.66% 6.63
10 4.39 ($\pm$ 1.43) 80% 7.05
Best-of-N Lag. $N=128$ 0 7.83 ($\pm$ 1.19) 33% 7.92
1 6.47 ($\pm$ 1.43) 68.33% 8.46
2.5 5.84 ($\pm$ 1.51) 74.33% 9.0
5 5.22 ($\pm$ 1.39) 79.3% 9.08
10 4.62 ($\pm$ 1.37) 80.67% 8.83
Beam Search Lag. $N=64$ 0 7.98 ($\pm$ 1.62) 36% 6.8
1 6.96 ($\pm$ 1.67) 76.33% 7.29
2.5 6.28 ($\pm$ 1.66) 79.33% 7.57
5 5.36 ($\pm$ 1.42) 83% 7.34
10 4.91 ($\pm$ 1.49) 84.67% 7.11
Beam Search Lag. $N=128$ 0 8.44 ($\pm$ 1.64) 32.67% 11.83
1 7.43 ($\pm$ 1.76) 77.3% 11.8
2.5 6.53 ($\pm$ 1.75) 80.33% 11.55
5 5.70 ($\pm$ 1.57) 83% 11.28
10 5.02 ($\pm$ 1.51) 83.67% 11.41
ARGS Lag. 0 5.67 ($\pm$ 1.45) 47% 68.23
1 3.39 ($\pm$ 1.6) 83.67% 72.4
2.5 2.1 ($\pm$ 1.73) 92.67% 75.5
5 1.72 ($\pm$ 1.96) 93.33% 79.28
10 0.11 ($\pm$ 1.59) 93.33% 80.34
HH-RLHF InferenceGuard. $N=128$ 9.49 ($\pm$ 2.16) 98.97% 11.47
Best-of-N Lag. $N=64$ 0 9.14 ($\pm$ 1.99) 95.33% 5.62
1 7.80 ($\pm$ 1.89) 95.98% 6.07
2.5 7.44 ($\pm$ 2.09) 96.53% 5.71
5 6.67 ($\pm$ 2.48) 97.10% 6.83
10 5.44 ($\pm$ 2.69) 97.24% 6.64
Best-of-N Lag. $N=128$ 0 9.42 ($\pm$ 2.01) 95.6% 8.01
1 8.32 ($\pm$ 1.90) 95.84% 8.7
2.5 7.85 ($\pm$ 2.14) 96.87% 8.24
5 6.97 ($\pm$ 2.54) 97.24% 8.31
10 5.47 ($\pm$ 2.88) 97.5% 8.79
Beam Search Lag. $N=64$ 0 9.14 ($\pm$ 1.99) 95.33% 7.61
1 8.87 ($\pm$ 2.29) 96.59% 6.92
2.5 8.47 ($\pm$ 2.37) 96.96% 7.04
5 7.88 ($\pm$ 2.20) 96.6% 8.02
10 6.80 ($\pm$ 2.42) 97.51% 9.07
Beam Search Lag. + Freq. $N=64$ 1 8.91 ($\pm$ 2.34) 96.8% 6.93
2.5 8.52 ($\pm$ 2.4) 97.01% 7.04
5 7.81 ($\pm$ 2.39) 97.27% 8.54
10 6.81 ($\pm$ 2.42) 97.51% 9.01
Beam Search Lag. $N=128$ 0 9.42 ($\pm$ 2.01) 95.6% 11.4
1 9.33 ($\pm$ 2.36) 96.4% 11.06
2.5 8.89 ($\pm$ 2.51) 97.44% 11.65
5 8.05 ($\pm$ 2.25) 97.54% 11.28
10 6.85 ($\pm$ 2.56) 97.77% 12.26
Beam Search Lag. + Freq. $N=128$ 1 9.40 ($\pm$ 2.31) 96.81% 11.29
2.5 8.92 ($\pm$ 2.33) 97.96% 11.64
5 8.05 ($\pm$ 2.54) 97.64% 11.32
10 6.85 ($\pm$ 2.55) 97.71% 12.53
ARGS Lag. 0 6.83 ($\pm$ 1.83) 96.2% 70.4
1 3.98 ($\pm$ 1.79) 97% 63.78
2.5 2.65 ($\pm$ 1.8) 96.4% 67.96
5 2.02 ($\pm$ 1.79 ) 97.54% 73.74
10 0.74 ($\pm$ 1.88 ) 97.99% 74.98
Table 7: Performance Comparison of InferenceGuard w.r.t. Alpaca-7B on Dataset PKU-SafeRLHF using Different $d$, $N$, and $K$ and fixed $D=128$
Method $K$ Average Reward Average Cost Safety Rate Inference Time (s)
Alpaca-7B InferenceGuard $N=128,d=16$ 64 6.6 ($\pm$ 2.5) -0.72 94.07% 28.45
32 7.14 ($\pm$ 2.75) -0.84 94.3% 27.15
16 7.64 ($\pm$ 2.85) -0.81 94.33% 25.66
InferenceGuard $N=128,d=32$ 64 5.98 ($\pm$ 2.5) -0.86 95.65% 14.70
32 6.39 ($\pm$ 2.7) -0.94 96.3% 13.38
16 6.66 ($\pm$ 2.74) -0.89 96.05% 14.22
InferenceGuard $N=128,d=64$ 64 5.5 ($\pm$ 2.46) -0.98 96.97% 7.82
32 5.71 ($\pm$ 2.5) -0.92 96.84% 7.85
16 5.82 ($\pm$ 2.61) -0.94 96.97% 5.84
InferenceGuard $N=256,d=16$ 128 6.83 ($\pm$ 2.5) -0.88 96.18% 42.24
64 7.56 ($\pm$ 2.81) -0.98 97.1% 37.77
32 7.73 ($\pm$ 2.93) -1 98.55% 36.92
InferenceGuard $N=256,d=32$ 128 6.19 ($\pm$ 2.51) -0.99 98.15% 22.66
64 6.67 ($\pm$ 2.73) -0.94 96.97% 22.38
32 6.99 ($\pm$ 2.90) -1.03 98.15% 22.77
InferenceGuard $N=256,d=64$ 128 5.82 ($\pm$ 2.6) -0.98 98.42% 5.82
64 5.92 ($\pm$ 2.63) -1.05 99.34% 9.89
32 6.08 ($\pm$ 2.72) -1.04 97.5% 11.28
InferenceGuard $N=64,d=16$ 32 6.76 ($\pm$ 2.46) -0.5 86.56% 16.44
16 7.28 ($\pm$ 2.59) -0.65 89.2% 15.32
8 7.45 ($\pm$ 2.69) -0.6 89.06% 15.29
InferenceGuard $N=64,d=32$ 32 5.95 ($\pm$ 2.42) -0.65 90.38% 12.09
16 6.48 ($\pm$ 2.5) -0.63 90.0% 11.58
8 6.64 ($\pm$ 2.63) -0.67 90.8% 11.73
InferenceGuard $N=64,d=64$ 32 5.67 ($\pm$ 2.41) -0.73 91.17% 6.17
16 5.79 ($\pm$ 2.46) -0.76 91.57% 6.23
8 5.81 ($\pm$ 2.45) -0.75 90.6% 3.93
Robustness Analysis of InferenceGuard
To evaluate the robustness of InferenceGuard, we vary the number of samples $N$, safety budget $d$, and selection width $K$. Table 7 reports results on the PKU-SafeRLHF dataset using Alpaca-7B. We observe that increasing $N$ and reducing $d$ generally improves task reward at the expense of inference latency, while higher $d$ leads to stronger safety enforcement. For instance, at $N=256$, InferenceGuard achieves a safety rate of up to $99.34\%$ while maintaining a competitive reward of 6.08.
Figure 4: Pareto curves show the safety-reward trade-offs for decoding methods on (1) Alpaca-7B with PKU-SafeRLHF and (2) Vicuna-7B with HEx-PHI. Each curve corresponds to a $\lambda$ or safety budget ablation, tracing the approximate Pareto front.
Finally, Figure 4 shows the approximate Pareto front for reward versus safety cost across decoding strategies. InferenceGuard consistently traces a more favorable Pareto curve, offering better safety-reward trade-offs compared to baseline methods on both Alpaca-7B (PKU-SafeRLHF) and Vicuna-7B (HEx-PHI).
D.4 Limitations and Latency Analysis
While InferenceGuard demonstrates robust safety improvements across diverse models and datasets, we also acknowledge several limitations that merit further discussion.
Table 8: Win-rate Percentage Comparison on PKU-SafeRLHF evaluated by ’Deepseek-r1-distill-qwen-32b’
Method Helpfulness Win Rate (%) Harmlessness Win Rate (%)
InferenceGuard with critic 72 76.8
InferenceGuard 66.8 76.6
BeamSearch-Saute (N=256) 68.6 75
BoN-Saute (N=500) 61.4 62.0
BoN-lagrange (N=500, $\lambda=5$) 67 60.6
Args-Lagrange 14.2 52.2
RE-Control-Lagrange 52 50.4
RE-Control 50.6 49.2
ARGS-Vanilla 51.2 47.6
Figure 5: Win, tie, and loss counts of alignment methods compared against responses generated by Alpaca-7B on the PKU-SafeRLHF dataset, using Deepseek-r1-distill-qwen-32b as the judge model.
Reliance on Cost Model Quality.
Our theoretical analysis suggests that if the cost model reliably distinguishes unsafe content, InferenceGuard can almost surely generate safe responses. In practice, cost models may exhibit bias, dataset artifacts, or false positives. To address this concern, we evaluate InferenceGuard under a stronger judge model (e.g., Deepseek-r1-distill-qwen-32b[124]) independently from the cost model (as shown in Table 8 and Figure 5), and it still outperforms baseline methods in safety under distribution shift, indicating robustness and reduced overfitting to the cost function.
Table 9: Average inference time per prompt on Beaver-7B evaluated on PKU-SafeRLHF. The total inference time per prompt is decomposed into generation time using vllm and evaluation time on the reward model, cost model or critic.
Method Num Samples Beam Depth Safety Rate (%) Generation Time (s) Eval Time (s) Inference Time (Gen + Eval) (s)
Best-of-N + Augmented Safety 32 N/A 88.14% 0.82 0.6 1.42
Best-of-N + Augmented Safety 64 N/A 91.17% 1.58 1.1 2.68
Beam Search + Augmented Safety 32 64 89.60% 1.34 0.6 1.94
Beam Search + Augmented Safety 64 64 90.07% 2.67 1.2 3.87
InferenceGuard 32 64 99.21% 1.32 0.7 2.02
InferenceGuard 64 64 99.60% 2.85 1.2 4.05
Latency Overhead.
Test-time Alignment methods incur additional computation due to multi-sample decoding and online evaluation. This is an inherent challenge for most alignment methods that aim to intervene during generation. Table 9 reports end-to-end inference time across models and decoding methods. Our latency is comparable to Beam Search, particularly on fine-tuned models such as Beaver-7B-v3. We regard this gap as an “alignment tax” [27] – a computational cost required to achieve higher safety guarantees, and it is manageable based on the preference between efficiency and safety. Importantly, InferenceGuard achieves significantly better safety than methods with similar runtime profiles, and our results highlight a trade-off between latency and safety, where InferenceGuard provides strong safety guarantees at a reasonable computational cost.
D.5 Critic Network and Training Process